OpenStudy (anonymous):

A Rectangular box with square base and open at the top is to have a capacity of 16,823 cu.cm. Find the height of the box that requires minimum amount of the material required. a. 14.12cm b. 16.14cm c. 12.13cm d. 10.36cm

OpenStudy (phi):

Perhaps the hardest part is coming up with the equations.

OpenStudy (hoblos):

let the side of base =x and height=h hx^2 = 16,823 so x = √(16823/h) surface area of our bos = S(x)= x^2 +4xh to find the minimum we calculate the first derivative S'(x) = 2x + 4h this is minimum for S'(x) =0 2x + 4h = 0 replace x by its value.. 2√(16823/h) + 4h = 0 4h = -2√(16823/h) square both sides 16h^2 = 4(16823/h) cross multiply h^3 = 4(16823)/16 so h = [(16823)/16]^(1/3) = 16.14cm

OpenStudy (hoblos):

sorry the last line.. h = [(16823)/4]^(1/3) = 16.14cm

OpenStudy (anonymous):

excellent, @hoblos

OpenStudy (hoblos):

thanks :D

OpenStudy (anonymous):

i reading hoblos answer i learn 1 or 2 w8.. im not finish =)

OpenStudy (anonymous):

amazing like a prof

OpenStudy (hoblos):

XD hope you understood that..

OpenStudy (phi):

Find the height of the box that requires minimum amount of the material required. That means find the least surface area (assumes all the sides and bottom are the same thickness). So you need the equation for the surface area (without the top). 4 sides and 1 bottom Then you need to replace the variable h (height) using the equation volume= h*x*x See hoblos for the details.